OpenStudy (anonymous):

Solve: −2(x − 8) = −10 −13 13 −3 3

OpenStudy (anonymous):

try to ..ans ur self we will check ....

OpenStudy (uri):

−2(x − 8) = −10 -2x +16 = -10 -2x =-10 -16 Can you do it now?

OpenStudy (anonymous):